Donald was killed in an accident while he was on the job in 2018. Darlene, Donald's wife, received several payments as a result of Donald's death. Review the payments below and then enter the amount to be included in Darlene's gross income in the table provided. Donald's employer paid Darlene an amount equal to Donald's three months' salary ($18,400), which is what the employer does for all widows and widowers of deceased employees. Donald had $6,800 in accrued salary that was paid to Darlene. Donald's employer had provided Donald with group term life insurance of $225,000, which was payable to his widow in a lump sum. Premiums on this policy totaling $21,000 had been included in Donald's gross income under § 79. Donald had purchased a life insurance policy (premiums totaled $156,000) that paid $376,000 in the event of accidental death. The proceeds were payable to Darlene, who elected to receive installment payments as an annuity of $35,500 each year for a 22-year period. She received her first installment this year.
Answer #1

Points to be considered in above ques:

1. Darlene received three months' salary from Donald's employer amounting to $18400

2. Darlene also received accrued salary of Donald amounting to $ 6800

3. Proceeds from Donald's Group term life insurance will be $0

4. Calculation of amount which Darlene will get from Donald's Life insurance Policy:

Installment payment as annuity is $35500 which will be deducted from capital recovery

a. Calculation of Recovery of Capital

(investment in the contract/expected return)* annuity amount

(376000/(35500*22))*35500

=(376000/781000)*35500

= 0.481*35500

=17090.9

b. Calculation of Amount which Darlene will receive from Insurance Policy :

$35500-17090.9= $18409.0

ANS: Darlene's Gross Income will be : 18400+6800+0+18409 = 43609